# Probing Loop Quantum Gravity via Kerr Black Hole and EHT Results

### v1

@@ Abstract
Recently, a study on optical properties and shadow of quantum Schwarzschild black hole appeared in [Ye et al., Phys. Lett. B 851, 138566, (2024)] for a fixed Barbero-Immirzi parameter $\gamma$. Following the same approach, we considered its rotating counterpart which is precisely a deformed Kerr metric in Loop Quantum Gravity. The deviation between the quantum-corrected Kerr and Kerr black holes has been investigated by the analysis of horizon structure and null geodesics by assuming a fixed value of $\gamma$. We have proved a theorem dealing with the location of unstable circular null orbits for all metrics of this kind by incorporating the convexity of effective potential of the Kerr black hole. The deviation between the shadows of the quantum-corrected and Kerr black holes has also been studied, and lastly the shadow analysis is incorporated in comparison with the EHT results for M87* and Sgr A* to precisely probe the quantity of deviation due to quantum correction. We have found that the quantum correction significantly reduces the extremal spin value and hence the size of the black hole as compared to Kerr black hole. Moreover, the unstable null orbits for quantum black hole are always smaller than the unstable null orbits for Kerr black hole. Lastly, we found that the quantum correction allows the deformed Kerr black hole to mimic Sgr A* with a higher probability than the Kerr black hole. However, the quantum-corrected Kerr black hole barely mimics M87*.
January 2, 2025
@@ I Introduction
In 2000, Falcke et al. [30] predicted the possibility of imaging the BHs especially Sgr A*, which got verified by the discovery of images of the supermassive BH M87* [31, 32] and Sgr A* [33, 34] by the Event Horizon Telescope (EHT) that has opened up new avenues in BH physics. These images comprise light rings and shadows of the BHs with nearly equal radii. A light ring is formed by the trapping of light and appears as a glowing image of radius greater than the event horizon. Whereas, the shadow is not a physical entity, but a dark 2D silhouette formed by the disappearing of photons from the sight of an observer at infinity. Since then, many studies have been accomplished to understand the shape and size of the BH shadow in various different models and frameworks, see Refs. [35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47]. The images of M87* and Sgr A* provided the data which has been useful in testing various gravity theories by determining constraints on the parameters associated with the theories, see Refs. [48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65]. In particular, Islam et al. [52] investigated LQG by using the EHT data for M87* and Sgr A*. They considered a rotating polymerized BHs in LQG which acts as Kerr BH asymptotically. Using the shadow analysis, the LQG parameter has been constrained and found that a significant part of parametric spaces for one and two horizon BHs is consistent with the EHT results for both M87* and Sgr A*. Moreover, the EHT results for Sgr A* also agree with the triple horizon BH, but not for the M87*. Afrin et al. [53] investigated the LQG for the results of M87* and Sgr A* by considering two rotating LQG-inspired BHs and found that the upper bound for LQG parameter obtained from the results of Sgr A* is more precise than the upper bound from M87*.
@@
Inspired by this kind of analysis, we consider the recently developed rotating counterpart [66, 67] of the quantum-corrected Schwarzschild BH [28], since the supermassive BHs are rotating in nature. Our major goal is to test the LQG effects under the influence of Barbero-Immirzi parameter, by comparing its shadow results with the EHT data and for Kerr BH. In particular, we investigate the deviation of the quantum-corrected Kerr BH from the classic Kerr BH via EHT results. The paper is organized as: In Sect. II, we present a brief overview of the quantum Oppenheimer-Snyder and Swiss-Cheese models and the obtained static BH metric. Further, we also present the rotating BH metric and discuss its horizon structure. In Sect. III, we employ the dynamical methods for the null geodesics, and effective potential and shadows are studied in comparison with result of Kerr BH. The Sect. IV comprises the constraints on the spin of BH parameters in comparison with EHT data. Lastly, in Sect. V, we present a brief conclusion and future prospects. Note that, we consider $G=\hbar=c=1$ in our calculations, unless otherwise mentioned. Moreover, the terms quantum BH, quantum Kerr BH, quantum-corrected BH, deformed BH, deformed Kerr BH, BH in LQG etc. refer to the same BH.
@@ II The Black Hole Metrics in Loop Quantum Gravity
$\text{d}s^{2}_{\text{qOS}}=-\left(1-F(r)\right)\text{d}t^{2}+\frac{\text{d}r^{2}}{1-G(r)}+r^{2}\text{d}\Omega_{2}^{2}.$ (3)
@@
In both cases, the dust interface $\tilde{r}=\tilde{r}_{0}$ is important region in APS spacetime which is described in the other spacetime with coordinates $(t(\tau),r(\tau),\theta,\phi)$. By smoothly matching the spacetimes on the interface so that the coordinates $(\tau,\tilde{r}_{0},\theta,\phi)\sim(t(\tau),r(\tau),\theta,\phi)$ generate the values of the functions $F(r)$ and $G(r)$. As a result, we get the deformed Schwarzschild metric given by [28]
@@
At the interface, the radius of the dust ball being $a(\tau)\tilde{r}_{0}$ is equal to the radius $r(\tau)$ in the other spacetime, therefore, taking the derivative on both sides and then squaring gives the value of $G(r(\tau))$. Along the radial geodesic, when $\dot{r}=0$, we have $G(r)=0$ which gives the lower bound of the radial coordinate, that is,
@@ IV Comparison With EHT Data
V Discussion
@@
By a smooth matching of APS and qOS spacetimes by identifying a set of suitable coordinate transformation, a deformed Schwarzschild BH in LQG is obtained with a relatively weak extra term of order $r^{-4}$ depending on Barbero-Immirzi parameter $\gamma$ [28]. During the collapse of the dust ball, corresponding to a lower bound of the radial coordinate, a lower bound on the mass of the quantum Kerr BH is obtained. Below which, there exists no horizon for the LQG BH, whereas, a two horizon system is obtained for all mass values above the minimum mass limit. Ye et al. [29] studied the shadows of static LQG BH to determine the effect of LQG on the Schwarzschild BH. They considered the fixed value of the Barbero-Immirzi parameter and hence the parameter $\alpha$. They performed the analysis of their findings and presented it in comparison with the results of Schwarzschild BH. Motivated by this, we considered the rotating counterpart of the deformed Schwarzschild BH in LQG and accomplished our analysis in comparison with the results for Kerr BH.

### v2

@@ Abstract
Recently, a study on shadow of quantum Schwarzschild black hole appeared in [Ye et al., Phys. Lett. B 851, 138566, (2024)] for a fixed value of Barbero-Immirzi parameter $\gamma$. Following this approach, we considered its rotating counterpart being a deformed Kerr metric in Loop Quantum Gravity and studied its deviation from Kerr black hole for a fixed value of $\gamma$. We proved a theorem describing the location of unstable circular null orbits for all metrics of this kind. The deviation between the shadows of the quantum Kerr and Kerr black holes has also been studied, and parameters are constrained by comparison with the EHT results for M87* and Sgr A* to precisely probe the quantity of deviation due to quantum correction. Lastly, we immersed the deformed Kerr black hole in an inhomogeneous plasma and studied its impact on the shadow size. We found that the unstable null orbits for the quantum black hole are always smaller than the unstable null orbits for Kerr black hole. The quantum correction allows the deformed Kerr black hole to mimic Sgr A* with a higher probability than the Kerr black hole. However, the quantum-corrected Kerr black hole barely mimics M87*. The plasma reduces the size of black hole shadow and the plasma parameter in the case II is more sensitive than that in case I.
February 23, 2025
@@ I Introduction
In 2000, Falcke et al. [30] predicted the possibility of imaging the BHs especially Sgr A*, which got verified by the discovery of images of the supermassive BH M87* [31, 32] and Sgr A* [33, 34] by the Event Horizon Telescope (EHT) that has opened up new avenues in BH physics. These images comprise light rings and shadows of the BHs with nearly equal radii. A light ring is formed by the trapping of light and appears as a glowing image of radius greater than the event horizon. Whereas, the shadow is not a physical entity, but a dark 2D silhouette formed by the disappearing of photons from the sight of an observer at infinity. Since then, many studies have been accomplished to understand the shape and size of the BH shadow in various different models and frameworks, see Refs. [35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47]. The images of M87* and Sgr A* provided the data which has been useful in testing various gravity theories by determining constraints on the parameters associated with the theories, see Refs. [48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65]. In particular, Islam et al. [52] investigated LQG by using the EHT data for M87* and Sgr A*. They considered a rotating polymerized BHs in LQG which acts as Kerr BH asymptotically. Using the shadow analysis, the LQG parameter has been constrained and found that a significant part of parametric spaces for one and two horizon BHs is consistent with the EHT results for both M87* and Sgr A*. Moreover, the EHT results for Sgr A* also agree with the triple horizon BH, but not for the M87*. Afrin et al. [53] investigated the LQG for the results of M87* and Sgr A* by considering two rotating LQG-inspired BHs and found that the upper bound for LQG parameter obtained from the results of Sgr A* is more precise than the upper bound from M87*. The astrophysical BHs are surrounded by plasma due to a huge gravity and in-falling matter [66]. This may certainly last an impact on the light propagation causing deviation in shadows. The study of light propagation in magnetized plasma with no pressure, is governed by the Hamiltonian derived by Breuer and Ehlers [67, 68]. Later, Perlick and Tsupko [69] derived the Hamiltonian formulation for a rather simpler case of a plasma without pressure and magnetization. Perlick et al. [70] also explored the shadow of static and spherically symmetric BHs immersed in a plasma medium. Recently, various BH solutions have been considered to investigate their shadows in the presence of plasma, see Refs. [71, 72, 73, 74, 75, 57, 76].
@@
Inspired by this kind of analysis, we consider the recently developed rotating counterpart [77, 78] of the quantum-corrected Schwarzschild BH [28], since the supermassive BHs are rotating in nature. Our major goal is to test the LQG effects under the influence of Barbero-Immirzi parameter, by comparing its shadow results with the EHT data and for Kerr BH. In particular, we investigate the deviation of the quantum-corrected Kerr BH from the classic Kerr BH via EHT results. We also study the impact of plasma on the shadow of deformed Kerr BH. The paper is organized as: In Sect. II, we present a brief overview of the quantum Oppenheimer-Snyder and Swiss-Cheese models and the obtained static BH metric. Further, we also present the rotating BH metric and discuss its horizon structure. In Sect. III, we employ the dynamical methods for the null geodesics, and effective potential and shadows are studied in comparison with result of Kerr BH. The Sect. IV comprises the constraints on the spin of BH parameters in comparison with EHT data. In Sect. V, we investigate the impact of plasma on shadows. Lastly, in Sect. VI, we present a brief conclusion and future prospects. Note that, we consider $G=\hbar=c=1$ in our calculations, unless otherwise mentioned. Moreover, the terms quantum BH, quantum Kerr BH, quantum-corrected BH, deformed BH, deformed Kerr BH, BH in LQG etc. refer to the same BH.
@@ II The Black Hole Metrics in Loop Quantum Gravity
$\text{d}s^{2}_{\text{qOS}}=-\left(1-F(r)\right)\text{d}t^{2}+\frac{\text{d}r^{2}}{1-H(r)}+r^{2}\text{d}\Omega_{2}^{2}.$ (3)
@@
In both cases, the dust interface $\tilde{r}=\tilde{r}_{0}$ is important region in APS spacetime which is described in the other spacetime with coordinates $(t(\tau),r(\tau),\theta,\phi)$. By smoothly matching the spacetimes on the interface so that the coordinates $(\tau,\tilde{r}_{0},\theta,\phi)\sim(t(\tau),r(\tau),\theta,\phi)$ generate the values of the functions $F(r)$ and $H(r)$. As a result, we get the deformed Schwarzschild metric given by [28]
@@
At the interface, the radius of the dust ball being $a(\tau)\tilde{r}_{0}$ is equal to the radius $r(\tau)$ in the other spacetime, therefore, taking the derivative on both sides and then squaring gives the value of $H(r(\tau))$. Along the radial geodesic, when $\dot{r}=0$, we have $H(r)=0$ which gives the lower bound of the radial coordinate, that is,
@@ IV Comparison With EHT Data
V Effect of Plasma on Shadow
We know that astrophysical BHs are generally surrounded by plasma. Moreover, excited states of matter at high temperatures with their constituent particles (positive and negative charges) may have significant impact on the LQG BH at Planck scales. This being an important aspect in theoretical physics, we investigate the influence of plasma on the appearance of quantum Kerr BH. We consider pressureless and non-magnetized plasma with distribution functions dependent on $r$ and $\theta$. We begin by considering the plasma electron frequency defined as
$\omega_{\text{p}}(r,\theta)=\frac{4\pi e^{2}}{m_{\text{e}}}N_{\text{e}}(r,\theta)$ (31)
that modifies the Hamiltonian describing the photon motion in a plasma medium given as [69]
$\mathcal{H}=\frac{1}{2}\left[g^{\mu\nu}p_{\mu}p_{\nu}+\omega^{2}_{\text{p}}(r,\theta)\right],$ (32)
where $e$, $m_{\text{e}}$ and $N_{\text{e}}$ are the charge, mass and number density of the electron, respectively. The plasma electron frequency is related to the refractive index $n(r,\theta)$ as
$n(r,\theta)=\sqrt{1-\frac{\omega^{2}_{\text{p}}(r,\theta)}{\omega^{2}(r,\theta)}},$ (33)
where $\omega(r,\theta)$ is the frequency of photon measured by a static observer outside the event horizon. The Hamilton’s equations are defined as
$\dot{x}^{\mu}=\partial_{\tau}x^{\mu}=\partial_{p_{\mu}}\mathcal{H},\qquad\dot{p}_{\mu}=\partial_{\tau}p_{\mu}=-\partial_{x^{\mu}}\mathcal{H}$ (34)
with $\tau$ being an affine parameter. Since, a plasma medium is dense and dispersive, therefore, the motion of photons is affected by the frequencies of electrons in the plasma. Therefore, the frequency of the propagating photon must be greater than the frequency of the electrons in plasma, that is, $\omega^{2}(r,\theta)\geq\omega^{2}_{\text{p}}(r,\theta)$. The observer is assumed static with the four-velocity in comoving coordinates given as $U^{\mu}(r,\theta)=\left(-g_{tt}(r,\theta)\right)^{-1/2}$ and the photon frequency is $\omega(r,\theta)=-p_{\mu}U^{\mu}(r,\theta)=-p_{t}U^{t}(r,\theta)$. The Planck’s relation connects the energy and the angular frequency as $E=\hbar\omega_{0}$ which gives $p_{t}=-\omega_{0}$ by setting the units for $\hbar=1$. Therefore, one obtains
$\omega(r,\theta)=\omega_{0}\left(-g_{tt}(r,\theta)\right)^{-\frac{1}{2}}.$ (35)
We obtain $t$ and $\phi$ components of geodesic equations corresponding to two constants of motion, energy $E$ and angular momentum $L$. For other two equations, we consider Hamilton-Jacobi equation in the form
$g^{\mu\nu}\partial_{x^{\mu}}\mathcal{S}\partial_{x^{\nu}}\mathcal{S}+\omega^{2}_{\text{p}}(r,\theta)=0$ (36)
with Jacobi action given by Eq. (12) with $m_{\text{p}}=0$ as the third constant of motion. For the given metric, the Hamilton-Jacobi equation
$\Delta(r)\left(\partial_{r}\mathcal{A}_{r}(r)\right)^{2}+\left(L^{2}\csc^{2}\theta-a^{2}E^{2}\right)\cos^{2}\theta+(L-aE)^{2}$
$+\left(\partial_{\theta}\mathcal{A}_{\theta}(\theta)\right)^{2}-\frac{\left(\left(r^{2}+a^{2}\right)E-aL\right)^{2}}{\Delta(r)}+\rho^{2}\omega_{\text{p}}^{2}(r,\theta)=0$ (37)
cannot be separated until we assume [69]
$\omega_{\text{p}}(r,\theta)=\frac{\sqrt{f_{r}(r)+f_{\theta}(\theta)}}{\rho},$ (38)
where, $f_{r}(r)$ and $f_{\theta}(\theta)$ are arbitrary functions that ensures the separability of the Eq. (37). Therefore, generating the Carter constant $\mathcal{Z}$ as fourth constant of motion, the null geodesic equations are of same form as Eqs. (13)-(16) but with the functions
$\mathcal{R}(r)$ $=$ $\left(aL-\left(r^{2}+a^{2}\right)E\right)^{2}$ (39)
$-\Delta(r)\left(f_{r}(r)+\mathcal{Z}+(aE-L)^{2}\right),$
$\Theta(\theta)$ $=$ $\mathcal{Z}+a^{2}E^{2}\cos^{2}\theta-L^{2}\cot^{2}\theta-f_{\theta}(\theta).$ (40)
The impact parameters $\xi$ and $\eta$ in presence of plasma become
$\xi(r_{\text{p}})$ $=\frac{1}{a\Delta^{\prime}(r_{\text{p}})}\Big{[}\left(r_{\text{p}}^{2}+a^{2}\right)\Delta^{\prime}(r_{\text{p}})-\Delta(r_{\text{p}})\Big{(}2r_{\text{p}}$
$+\sqrt{4r_{\text{p}}^{2}-f_{r}^{\prime}(r_{\text{p}})\Delta^{\prime}(r_{\text{p}})}\Big{)}\Big{]},$ (41)
$\eta(r_{\text{p}})$ $=\frac{1}{a^{2}\Delta^{\prime 2}(r_{\text{p}})}\Big{[}8r_{\text{p}}^{2}\Delta(r_{\text{p}})\big{[}a^{2}-\Delta(r_{\text{p}})\big{]}-\big{[}r_{\text{p}}^{4}$
$+a^{2}f_{r}(r_{\text{p}})\big{]}\Delta^{\prime 2}(r_{\text{p}})+\big{[}4r_{\text{p}}^{3}-\big{(}a^{2}$
$-\Delta(r_{\text{p}})\big{)}f_{r}^{\prime}(r_{\text{p}})\big{]}\Delta(r_{\text{p}})\Delta^{\prime}(r_{\text{p}})+2r_{\text{p}}\Delta(r_{\text{p}})\big{[}2\big{(}a^{2}$
$-\Delta(r_{\text{p}})\big{)}+r_{\text{p}}\Delta^{\prime}(r_{\text{p}})\big{]}\sqrt{4r_{\text{p}}^{2}-f_{r}^{\prime}(r_{\text{p}})\Delta^{\prime}(r_{\text{p}})}\Big{]}.$ (42)
Now considering two cases for the functions $f_{r}(r)$ and $f_{\theta}(\theta)$, we will construct celestial coordinates to investigate shadows. The case I corresponds to the assumption $f_{r}(r)=\omega_{\text{c}}^{2}\sqrt{M^{3}r}$ and $f_{\theta}(\theta)=0$, where, $\omega_{c}$ has the dimensions of frequency. The celestial coordinates for this case are of the same form as in Eqs. (24) and (25) with their respective impact parameters given in Eqs. (41) and (42). For the equatorial observer, the celestial coordinates reduce to the same form as in Eq. (26). The case II corresponds to the assumption $f_{r}(r)=0$ and $f_{\theta}(\theta)=\omega_{\text{c}}^{2}M^{2}\left(1+2\sin^{2}\theta\right)$ with the same parameter $\omega_{c}$. The horizontal component of the celestial coordinates for this case is of the same form as in Eq. (24), however, the vertical component is modified. For an observer at $\left(\infty,\theta_{0}\right)$, it becomes
Figure 5: Influence of $\omega_{\text{c}}$ in case I (upper panel) and case II (lower panel) on shadows of quantum Kerr BH for a fixed value of $a$.
$Y(r_{\text{p}})$ $=$ ±$\pm\Big{[}\eta(r_{\text{p}})+a^{2}\cos^{2}\theta_{0}-\xi^{2}(r_{\text{p}})\cot^{2}\theta_{0}$ (43)
$-\omega_{\text{c}}^{2}M^{2}\left(1+2\sin^{2}\theta_{0}\right)\Big{]}^{\frac{1}{2}}$
and for the equatorial observer, it reduces to
$Y(r_{\text{p}})=\pm\sqrt{\eta(r_{\text{p}})-3\omega_{\text{c}}^{2}M^{2}}.$ (44)
For both of above mentioned cases, we have plotted the shadows in Fig. 5 for a few values of plasma parameter $\omega_{\text{c}}$ and keeping a fixed spin of the BH. The spin deviates the shadow from a pure circular loop causing a distortion. While, the plasma parameter $\omega_{\text{c}}$ reduces the size of the shadow. It can also be seen that $\omega_{\text{c}}$ in case II is more significant and has higher sensitivity as compared to the case I. The variation in the shadow size for case II is greater than the variation observed in case I.
VI Discussion
@@
By a smooth matching of APS and qOS spacetimes by identifying a set of suitable coordinate transformation, a deformed Schwarzschild BH in LQG is obtained with a relatively weak extra term of order $r^{-4}$ depending on Barbero-Immirzi parameter $\gamma$ [28]. During the collapse of the dust ball, corresponding to a lower bound of the radial coordinate, a lower bound on the mass of the quantum Kerr BH is obtained. Below which, there exists no horizon for the LQG BH, whereas, a two horizon system is obtained for all mass values above the minimum mass limit. Ye et al. [29] studied the shadows of static LQG BH to determine the effect of LQG on the Schwarzschild BH. They considered the fixed value of the Barbero-Immirzi parameter and hence the parameter $\alpha$. They presented their findings in comparison with the results of Schwarzschild BH. Motivated by this, we considered the rotating counterpart of the deformed Schwarzschild BH in LQG and accomplished our analysis in comparison with the results for Kerr BH.
@@
The plasma surrounding the LQG BH has a great impact on the light propagation. The shadow size is reduced by increasing the value of $\omega_{\text{c}}$ in both cases. However, the quantity of variation in shadow size reveals that $\omega_{\text{c}}$ in case II is more sensitive and has greater impact than that in case I.